OpenStudy (anonymous):

Find the distance between the points (-7, 5) and (4, 1).

OpenStudy (anonymous):

distance from -7 to 4 is 11 distance form 5 to 1 is 4 \[\sqrt{11^2+4^2}=\sqrt{121+16}=\sqrt{137}\]